# DeFiScope: Detecting Various DeFi Price Manipulations with LLM Reasoning

### v1

@@ 4.1. LLM Fine-tuning
For fine-tuning techniques, we chose OpenAI’s fine-tuning paradigm (OpenAi, 2024a) instead of supervised fine-tuning (SFT) (Chung et al., 2024) and its parameter-efficient version, LoRA (Hu et al., 2021), because a very small set of training data is required for the former while much more data points are needed for the latter (Ma et al., 2025). Moreover, OpenAI’s GPT family models demonstrate state-of-the-art reasoning capabilities on common benchmarks (Analysis, 2024; Sun et al., 2024a), which provides a suitable foundation for our fine-tuning. Accordingly, DeFiScope has enhanced the GPT-3.5-Turbo and GPT-4o models with data synthesized using the commonly used price calculation model, i.e., CPMM as illustrated in §2.2, along with on-chain data to fine-tune them.
@@
On-Chain Data Simulation. We leverage the fuzz testing method in Foundry (Github, 2024), an off-the-shelf toolkit for Ethereum application development, to simulate on-chain data. To avoid data leakage and generate a substantial volume of transactions satisfying the CPMM, we select the Uniswap V2:BTC20 (Etherscan, 2024b) liquidity pool as our target. We randomly generate inputs, namely integers ranging from $10^{20}$ to $10^{21}$ — 100 Ether to 1000 Ether, for the swap operations which are simulated on a forked blockchain of block height 17,949,214. Specifically, to include the data of inflating the price of tokens, we craft particular operations. To begin with, we record the balance of WETH and BTC20 in the liquidity pool denoted as $bal_{WETH}$ and $bal_{BTC20}$ respectively. Then we trigger swapExactTokensForTokens in contract UniswapV2Router02 (Etherscan, 2024c) to swap a amount of BTC20 for WETH, and record the latest balance of WETH, $bal_{WETH}^{{}^{\prime}}$, and that of BTC20, $bal_{BTC20}^{{}^{\prime}}$. Finally, we obtained the tokens’ balance change as a pair ($bal_{WETH}-bal_{WETH}^{{}^{\prime}},bal_{BTC20}-bal_{BTC20}^{{}^{\prime}}$). In terms of deflating the price of tokens, we swap a amount of WETH for BTC20 instead, with similar subsequent operations. Finally, we build a fine-tuning database comprising 500 pairs for price inflating and 500 pairs for price deflating, respectively. Despite the only use of CPMM-based DeFi protocols, our evaluation results in §7 demonstrates a significant gain in term of price manipulation attack detection for DeFi protocols using custom price models.
CoT-style Fine-tuning.
@@
Figure 4. The prompt template used in fine-tuning the LLM.
@@
Figure 4 demonstrates the prompt template used in the fine-tuning. We construct a CoT-style fine-tuning prompt that integrates both on-chain data and the price context. Above the dashed line is the first instruction, which requires the LLM to extract the price calculation model from the provided code. {code} is the placeholder for the code snippet of price calculation functions. Below the dashed line, we guide the LLM to evaluate the credibility of four statements based on the price model extracted from step 1 and the tokens’ balance change. We demand that the LLM expresses the credibility of a statement using integers ranging from 1 to 10. Compared to merely responding with a simple “Yes” or “No,” this scoring method also indicates the confidence level of the responses, which can assist us in selecting the answers in which the LLM is more confident.
@@
In the data part of the template, {$value_{0}$} and {$value_{1}$} are the placeholders for the first and second values in a price change pair, which are sampled from the fine-tuning dataset, respectively. {direction of change} can be either “increases” or “decreases.” Specifically, if {$value_{i}$} is greater than 0, {direction of change} is “increases”; conversely, if {$value_{i}$} is less than 0, {direction of change} is “decreases.” For the answer part, {score} is the score placeholder, which can be an integer between 1 and 10.
@@
Following OpenAI’s fine-tuning guideline (OpenAi, 2024a) that recommends using 50 to 100 training examples, we randomly sample 96 non-repetitive (in current and previous training sets) data from the fine-tuning database, and allocate 83% of the samples for training and 17% for validation. Subsequently, we insert data from the training and validation sets into the prompt template. To obtain the desired response for each prompt, we firstly ask LLMs to generate raw response, including the analysis of price model and scores of statements, for the given prompt. Next, we manually verify the correctness of the responses, of which the correct responses are stored and the wrong responses will also be corrected. Particularly, we simply swap the scores of two opposite statements to correct the error in them. Through this process, we construct a ground truth about prompts and its responses for fine-tuning the LLM model.
@@
During the fine-tuning process, training hyperparameters were automatically configured by the OpenAI fine-tuning API used. In particular, once the model achieved 100% accuracy on the validation set, we terminated the training to avoid overfitting. The entire fine-tuning process is cost-efficient, consuming a total of around 1 million training tokens for each model. The detailed costs across different models will be introduced in §7.2.
@@
Figure 5. The simplified Type-I prompt and its response for the motivating example during the inference process.
4.2. Inference with the Fine-tuned LLM
@@
In this section, we first illustrate the general inference process using the motivating example. Nevertheless, there could be closed source DeFi protocols that dissatisfy steps ⑤⑥ so that we design a customized inference process to deal with them.
@@
General Inference Process. To assist LLMs in extracting and analyzing the price model, our inference prompt provides relevant code for LLMs to score the trend of price changes and guides them in inferring the token price changes. Figure 5 illustrates a simplified version of the prompt used and the response produced by our fine-tuned LLM for inferring price changes of the motivating example in §2.3. We call this Type-I prompt, used for the typical cases with code input retrieved from steps ⑤⑥. It is different from the fine-tuning prompt in the {statement} and {change_description} parts. In the fine-tuning prompt shown in Figure 4, both parts are fixed, while they are dynamically generated during inference using two formats: (i) “The price of {token_name} in {contract_name} {direction_of_change} after change” for the placeholder {statement}; and (ii) “The balance of {token_name} in {contract_name} {direction_of_change} by {change_value}” or “The total supply of {token_name}{direction_of_change} by {change_value}” for the placeholder {change_description}. However, to fill them in Type-I prompt, DeFiScope generates a pair of statements for each token, i.e., one regarding the increase in token price and another regarding the decrease in token price.
@@
DeFiScope asks the fine-tuned LLM to locate the price calculation model from the input code and evaluate the credibility of the generated statements. From the motivating example’s response shown in the right-hand section of Figure 5, the LLM initially extracts the code of price calculation-related functions, followed by an high-level summary. In this example, it accurately identifies the underlying price model (see eq. 1) — the price of sUSDe is determined by the median of multiple prices. With this knowledge learned, the LLM could credit two opposite statements given the tokens balance changes and yields the correct answer with high confidence.
Figure 6. The Type-II prompt for closed-source liquidity pools.
@@
Customized Inference Process. Although the majority of DeFi applications are open source to gain users’ trust, some remain closed source, making our Type-I prompt inapplicable. To address this, we developed a Type-II prompt template, as shown in Figure 6, to infer the trend of price changes in closed source two-token liquidity pools. Our observation is that the majority of two-token liquidity pools pools use the CPMM as their underlying price model. Therefore, the primary distinction between the Type-I prompt and the Type-II prompt lies in replacing the first instruction with a description of the liquidity pool, informing the LLM that the pool’s price model aligns with CPMM. It is worth noting that the liquidity pool is automatically identified during transaction analysis, which will be introduced in §5. Due to page limitation, we include a case study of using the Type-II prompt for inference in Appendix A.
5. DeFi Operations
@@
To bridge this gap, we first model token transfer actions using a directed graph (§5.1), and then recover high-level DeFi operations from it (§5.2). Since our detection is based on a single transaction, it should be noted that all described operations are derived from one raw transaction, and we do not consider the DeFi operations expressed by the combination of multiple raw transactions. Based on our study of the top-10 high-value DeFi applications across all three categories mentioned in §2.1 (the full list is available in Appendix B) with active transactions in each category, due to the susceptibility to front-running across multiple transactions and the atomicity of transactions ensuring complete execution of operations, only a very few DeFi operations span multiple transactions.
5.1. Transfer Graph Construction
We define the Transfer Graph (TG) (Definition 5.2), a directed graph where the edges represent transfer actions (Definition 5.1) and the vertices represent related accounts, to model transfer actions within each user invocation.
Definition 0 (Transfer).
@@ 5.1. Transfer Graph Construction
Definition 0 (Transfer Graph).
@@
A Transfer Graph (TG) is a tuple ($\mathcal{A}$$\mathcal{E}$), where $\mathcal{A}$ is the set of all accounts (including EOAs, CAs and $\emptyset$) involved in a user invocation, $\mathcal{E}$ is the set of directed edges, i.e., =$\mathcal{E}=\{E_{1},...,E_{m}\}\subseteq\mathcal{A}\times\mathcal{A}$, where each $E_{i}:=\langle j,T_{k}\rangle$, j is the time index of $T_{k}$, $T_{k}\in\mathcal{T}$, $\mathcal{T}$ is the set of all transfer actions involved in the user invocation, i.e., $\mathcal{T}=\{T_{1},...,T_{n}\}$, where $T_{i}.s,T_{i}.r\in\mathcal{A}$ for each $T_{i}$.
@@
According to our categorization, a transfer action can be one of three types: transferring token, burning token, and minting token. All transfer actions can be expressed as “Sender transfers amount of token to Receiver.” In a transferring token action, all accounts involved must be either EOAs or CAs, and must not be a zero address2220x0000000000000000000000000000000000000000 or a dead address3330x000000000000000000000000000000000000dEaD (we uniformly denote these two special addresses by $\emptyset$). Meanwhile, the receiver in a burning token action and the sender in a minting token action must be $\emptyset$.
@@
Figure 7, as demonstrated in step ①, illustrates the construction of a TG from the raw transaction of a user invocation. This user invocation includes six contract accounts and a collection of user-controlled accounts $UC$, which includes EOAs and CAs, along with seven transferring token actions. The Sender and Receiver of a transfer action are connected by a directed edge, from the Sender to the Receiver, with a time index to indicate the order of occurrence. $T_{1}$, from one of the user-controlled accounts in $UC$ to $CA_{1}$, is the first transfer action in this user invocation. Before $CA_{1}$ transfers tokens to $CA_{2}$ through $T_{4}$, user-controlled accounts initiate two transfers to $CA_{4}$ and $CA_{5}$ respectively, resulting in $T_{4}$ having a larger time index compared to $T_{2}$ and $T_{3}$. Similarly, since $T_{6}$ occurs between $T_{5}$ (with a time index of 5) and $T_{7}$ (with a time index of 7), its time index is set to 6.
Compared to existing methods that model transfer actions in raw transactions utilized for price manipulation detection, such as the Cash Flow Tree (CFT) constructed in DeFiRanger (Wu et al., 2023), the novelty of the Transfer Graph (TG) lies in its unique modeling approach. The CFT models invocation and transfer actions within raw transactions, where the edges represent invocation relationships and the vertices represent a called function or a transfer action. In contrast, the TG models transfer actions in each user invocation within raw transactions, thereby offering greater granularity. Furthermore, the TG employs a directed graph rather than a tree structure for modeling and retains temporal relationships between different transfer actions, which enables it to cover more complicated DeFi operations and accelerate the speed. The proposed transfer graph are different from those fund flow graphs in prior works, such as  (McLaughlin et al., 2023). TG differs from the existing works in terms of granularity and design purpose. First, their directed graphs are built on transactions, while TG is defined on user invocations where a transaction can have multiple user invocations. Second, their directed graphs are to analyze transaction-related fund flow behaviors like arbitrage. In contrast, TG focuses on extracting high-level DeFi operations with the fund flow among user invocations.
@@
Figure 7. An illustrative example showing the workflow of recovering DeFi operations. $UC$: a collection of user-controlled accounts, including EOAs and CAs; $CA_{i}$: the $i$-th contract account; $T_{i}$: the $i$-th transferring token action.
5.2. DeFi Operation Inference
Based on our in-depth analysis of the top-30 DeFi apps shown in Table 7 (see Appendix B), we identify six DeFi operations that need to be recovered: 1) Swap, primarily from DEXs such as Uniswap; 2) Deposit, defined by yield-farming and lending apps such as AAVE and Pendle; 3) Withdraw, mainly from yield-farming apps; 4) Borrow, based on lending protocols like Compound; 5) Stake; and 6) Claim, both predominantly from yield-farming apps like Convex Finance.
Swap involves a user exchanging one token ($Token_{in}$) for another ($Token_{out}$) from a liquidity pool. Relevant contract accounts, excluding user-controlled ones, are noted as liquidity pools for price trend analysis when source code is unavailable. In a Deposit, a user transfers a token ($Token_{deposit}$) to a yield-farming or lending protocol and receives a proof token ($Token_{proof}$) through minting. Withdraw occurs when a user retrieves tokens ($Token_{withdraw}$) from a protocol by burning a proof token ($Token_{proof}$). Borrow refers to a DeFi operation where a borrower receives a token ($Token_{borrow}$) and incurs a debt evidenced by another token ($Token_{debt}$), issued through minting. Stake happens in yield-farming protocols like Convex Finance (Etherscan, 2024a), where users deposit tokens and can earn rewards without receiving a minted proof token. Claim enables users to retrieve staked tokens ($Token_{claim}$) and bonuses without burning any tokens, unlike the Withdraw operation which requires burning tokens to retrieve assets.
An Operation Recovering Example. To recover DeFi operations from the TG, we design and employ a search algorithm based on directed graphs for each operation. Since the recovery algorithm of DeFi operations is not the central focus of this work, we use the recovery of the Swap operation as an illustrative example here.
Figure 7 illustrates that, to recover a Swap operation, we use the depth-first search algorithm to identify cycles that both start and end at user-controlled accounts. These cycles must satisfy three constraints: (i) the transfer action of each edge must be transferring token; (ii) the time index of each edge must be monotonically increasing; and (iii) the token involved in the first transfer action must differ from the token in the last transfer action. Therefore, from the well-constructed TG, we can recover the Swap operation expressed as $UC\rightarrow T_{1}\rightarrow CA_{1}\rightarrow T_{4}\rightarrow CA_{2}\rightarrow T_{5}\rightarrow CA_{3}\rightarrow T_{7}\rightarrow UC$, and we label $CA_{1}$, $CA_{2}$, and $CA_{3}$ as liquidity pools.
Table 1. Price change information-directed attack patterns, systematically mined from all types of DeFi applications (see §2.1).

### v2

@@ IV-A LLM Fine-tuning
For fine-tuning techniques, we by default use OpenAI’s fine-tuning paradigm [22] instead of supervised fine-tuning (SFT) [23, 24], because the former requires only a small set of training data, whereas the latter generally requires more data points [25]. Nonetheless, we also demonstrate the generalizability of DeFiScope’s fine-tuning with open-source SFT models; see details in §VIII. In the default setting, DeFiScope enhances the GPT-3.5-Turbo and GPT-4o models with data synthesized using the CPMM price model as illustrated in §II-A, along with on-chain data to fine-tune them.
Our fine-tuning process comprises two components: (a) on-chain data simulation, which generates realistic token balance-change pairs from transactions on a forked blockchain network; and (b) chain-of-thought (CoT)-style fine-tuning, where these simulated data are integrated into CoT-style prompts that are then used to fine-tune the LLM, enabling it to perform stepwise reasoning.
@@
On-Chain Data Simulation. We leverage the fuzz testing method in Foundry [6], an off-the-shelf toolkit for Ethereum application development, to simulate on-chain data. To avoid data leakage and generate a substantial volume of transactions satisfying the CPMM, we select the Uniswap V2:BTC20 [26] liquidity pool as our target. We randomly generate inputs, namely integers ranging from $10^{20}$ to $10^{21}$ — 100 Ether to 1000 Ether, for the swap operations which are simulated on a forked blockchain of block height 17,949,214. Specifically, to include the data of inflating the price of tokens, we craft particular operations. To begin with, we record the balance of WETH and BTC20 in the liquidity pool denoted as $bal_{WETH}$ and $bal_{BTC20}$ respectively. Then we trigger swapExactTokensForTokens in contract UniswapV2Router02 [27] to swap a amount of BTC20 for WETH, and record the latest balance of WETH, $bal_{WETH}^{{}^{\prime}}$, and that of BTC20, $bal_{BTC20}^{{}^{\prime}}$. Finally, we obtained the tokens’ balance change as a pair ($bal_{WETH}-bal_{WETH}^{{}^{\prime}},bal_{BTC20}-bal_{BTC20}^{{}^{\prime}}$). In terms of deflating the price of tokens, we swap a amount of WETH for BTC20 instead, with similar subsequent operations. Finally, we build a fine-tuning database with 500 pairs each for price inflating and deflating. Despite the only use of CPMM-based DeFi protocols, our evaluation results in §VII demonstrates a significant gain in term of price manipulation attack detection for DeFi protocols using custom price models.
@@
Figure 2: The simplified prompt template used in fine-tuning the LLM. The full version can be found in our supplementary material [28].
@@
CoT-style Fine-tuning. Fig. 2 demonstrates the prompt template used in the fine-tuning. We carefully construct a CoT-style fine-tuning prompt for guiding the LLM toward step-by-step reasoning, which integrates both simulated on-chain data collected in the previous step and the price context. Above the dashed line is the first instruction, which requires the LLM to extract the price calculation model from the provided code. {code} is the placeholder for the code snippet of price calculation functions. Below the dashed line, we guide the LLM to evaluate the credibility of four statements based on the price model extracted from step 1 and the tokens’ balance change. We demand that the LLM expresses the credibility of a statement using integers ranging from 1 to 10. Compared to merely responding with a simple “Yes” or “No”, this scoring method also indicates the confidence level of the responses, which can assist us in selecting the answers in which the LLM is more confident. By offering a finer-grained evaluation scale, it reveals when the model is uncertain—if two contradictory statements receive the same confidence score, we know it cannot distinguish which is more trustworthy and therefore discard those low-confidence results.
@@
In the data part of the template, $value_{0}$} and $value_{1}$} are the placeholders for the first and second values in a price change pair, which are sampled from the fine-tuning dataset, respectively. {direction of change} can be either “increases” or “decreases.” Specifically, if $value_{i}$} is greater than 0, {direction of change} is “increases”; conversely, if $value_{i}$} is less than 0, {direction of change} is “decreases.” For the answer part, {score} is the score placeholder, an integer between 1 and 10.
@@
Following OpenAI’s fine-tuning guideline [22] that recommends using 50 to 100 training examples, we randomly sample 96 non-repetitive (in current and previous training sets) data from the fine-tuning database, and allocate 83% of the samples for training and 17% for validation. Subsequently, we insert data from the training and validation sets into the prompt template. To obtain the desired response for each prompt, we firstly ask LLMs to generate raw response, including the analysis of price model and scores of statements, for the given prompt. Next, we manually verify the correctness of the responses, of which the correct responses are stored and the wrong responses will also be corrected. Particularly, we simply swap the scores of two opposite statements to correct the error in them. Through this process, we construct a ground truth about prompts and its responses for fine-tuning.
@@
During the fine-tuning process, training hyperparameters were automatically configured by the OpenAI fine-tuning API (for SFT fine-tuning, hyperparameters were manually set using common practice; see §VIII). In particular, once the model achieved 100% accuracy on the validation set, we terminated the training to avoid overfitting. The entire fine-tuning process is cost-efficient, consuming a total of around 1 million training tokens for each model. The detailed costs across different models will be introduced in §VII-B.
@@
Figure 3: The simplified Type-I prompt and its response for the motivating example during the inference process.
IV-B Inference with the Fine-tuned LLM
@@
We first illustrate the general inference process using the motivating example. Nevertheless, there could be closed source DeFi protocols that dissatisfy steps ⑤⑥ so that we design a customized inference process to deal with them.
@@
General Inference Process. To assist LLMs in extracting and analyzing the price model, our inference prompt provides relevant code for LLMs to score the trend of price changes and guides them in inferring the token price changes. Fig. 3 illustrates a simplified version of the prompt used and the response produced by our fine-tuned LLM for inferring price changes of the motivating example in §II-B. We call this Type-I prompt, used for the typical cases with code input retrieved from steps ⑤⑥. It is different from the fine-tuning prompt in the <statement> and <change_description> parts. In the fine-tuning prompt shown in Fig. 2, both parts are fixed, while they are dynamically generated during inference using two formats: (i) “The price of {token_name} in {contract_name} {direction_of_change} after change” for the part <statement>, and (ii) “The balance of {token_name} in {contract_name} {direction_of_change} by {change_value}” or “The total supply of {token_name} {direction_of_change} by {change_value}” for the part <change_description>. Specifically, to fill them in Type-I prompt, DeFiScope generates a pair of statements for each token, i.e., one regarding the increase in token price and another regarding the decrease in token price.
@@
DeFiScope asks the fine-tuned LLM to locate the price calculation model from the input code and evaluate the credibility of the generated statements. From the motivating example’s response shown in the right-hand section of Fig. 3, the LLM first extracts the code of price calculation-related functions, followed by a high-level summary. In this example, it accurately identifies the underlying price model (see Equation 1): the price of sUSDe is determined by the median of multiple prices. With this knowledge, the LLM can evaluate two opposing statements based on the changes in token balances and provide the correct answer with high confidence, i.e., a higher score.
@@
Customized Inference Process. Although the majority of DeFi applications are open source to gain users’ trust, some remain closed source, making our Type-I prompt inapplicable. To address this, we developed a Type-II prompt template to infer the trend of price changes in closed source two-token liquidity pools. Our observation is that the majority of two-token liquidity pools use the CPMM as their underlying price model. Therefore, the primary distinction between the Type-I prompt and the Type-II prompt lies in replacing the Instruction 1 with a description of the liquidity pool, informing the LLM that the pool’s price model aligns with CPMM. It is worth noting that the liquidity pool is automatically identified during transaction analysis, which will be introduced in §V. To evaluate its effectiveness, we applied Type-II on a test dataset [29] of 100 price changes on CPMM, and it reached 97% accuracy, close to the 99% accuracy of Type-I. This demonstrates that, if applied to previously identified two-token liquidity pools, which commonly adopt CPMM, Type-II is substantially effective comparable to Type-I, while also adapting to closed-source pools. More details about the Type-II prompt and a case study of using it are available in [28].
V DeFi Operations
@@
To bridge this gap, we first model token transfer actions using a directed graph (§V-A), and then recover high-level DeFi operations from it (§V-B). Since our detection is based on a single transaction, it should be noted that all described operations are derived from one raw transaction, and we do not consider the DeFi operations expressed by the combination of multiple raw transactions. Based on our study of the top-10 high-value DeFi applications (the full list could be found in [28]) across three categories — Decentralized EXchange (DEX), Lending, Yield Farming, with active transactions in each category, due to the susceptibility to front-running across multiple transactions and the atomicity of transactions ensuring complete execution of operations, only a very few DeFi operations span multiple transactions.
V-A Transfer Graph Construction
We define the Transfer Graph (TG) (Definition 2), a directed graph where the edges represent transfer actions (Definition 1) and the vertices represent related accounts, to model transfer actions within each user invocation.
Definition 1 (Transfer)
@@ V-A Transfer Graph Construction
Definition 2 (Transfer Graph)
@@
A Transfer Graph (TG) is a tuple ($\mathcal{A}$$\mathcal{E}$), where $\mathcal{A}$ is the set of all accounts (including EOAs, CAs and $\emptyset$) involved in a user invocation, $\mathcal{E}$ is the set of directed edges, i.e., $\mathcal{E}=\{E_{1},...,E_{m}\}\subseteq\mathcal{A}\times\mathcal{A}$, where each $E_{i}:=\langle j,T_{k}\rangle$, j is the time index of $T_{k}$, $_{k}\in\mathcal{T}, 𝒯\mathcal{T}$ is the set of all transfer actions involved in the user invocation, i.e., $\mathcal{T}=\{T_{1},...,T_{n}\}$, where $T_{i}.s,T_{i}.r\in\mathcal{A}$ for each $T_{i}$.
@@
According to our categorization, a transfer action can be one of three types: transferring, burning, and minting token. All transfer actions can be expressed as “Sender transfers amount of token to Receiver.” In the transferring token, all accounts involved must be either EOAs or CAs, and must not be a zero address or a dead address (we uniformly denote these two special addresses by $\emptyset$). Meanwhile, the receiver in burning token and the sender in minting token actions must be $\emptyset$.
@@
Fig. 4, as demonstrated in step ①, illustrates the construction of a TG from the raw transaction of a user invocation. This user invocation includes six contract accounts and a collection of user-controlled accounts $UC$, which includes EOAs and CAs, along with seven transferring token actions. The Sender and Receiver of a transfer action are connected by a directed edge, from the Sender to the Receiver, with a time index to indicate the order of occurrence. $T_{1}$, from one of the user-controlled accounts in $UC$ to $CA_{1}$, is the first transfer action in this user invocation. Before $CA_{1}$ transfers tokens to $CA_{2}$ through $T_{4}$, user-controlled accounts initiate two transfers to $CA_{4}$ and $CA_{5}$ respectively, resulting in $T_{4}$ having a larger time index compared to $T_{2}$ and $T_{3}$. Similarly, since $T_{6}$ occurs between $T_{5}$ (with a time index of 5) and $T_{7}$ (with a time index of 7), its time index is set to 6.
Compared to the CFT approach used in DeFiRanger [4], which models invocation and transfer actions within raw transactions, our TG provides finer granularity. TG represents transfers as a directed graph while preserving their temporal order, enabling it to capture more complex operations more effectively. In our evaluation on 1,000 real-world Ethereum transactions, TG achieved a TPR ($\frac{\#TP}{\#TP+\#FN}$) of 0.912, significantly higher than CFT’s 0.559, while maintaining comparable precision ($\frac{\#TP}{\#TP+\#FP}$), i.e., 0.984 vs. 0.991. TG further demonstrated substantial advantages in recovering complex swap operations involving multiple relayers, achieving an 89.3% improvement over CFT. More details about the experimental setup and analysis are available in  [28]. Compared to prior fund-flow graph approaches, such as  [30], TG differs in two key aspects: (i) it is constructed at the level of user invocations rather than transactions, which may contain multiple invocations, and (ii) it is designed for high-level DeFi operation recovery, whereas those primarily target behavior analysis (e.g., arbitrage).
@@
Figure 4: An illustrative example showing the workflow of recovering DeFi operations. UCUC: a collection of user-controlled accounts, including EOAs and CAs; $CA_{i}$: the $i$-th contract account; $T_{i}$: the $i$-th transferring token action.
V-B DeFi Operation Inference
TABLE I: Price change information-directed attack patterns.